# PGSR: Planar-based Gaussian Splatting for Efficient and High-Fidelity Surface Reconstruction

### v1

@@ V-B Reconstruction
We compared our method, PGSR, with current state-of-the-art neural surface reconstruction methods including NeuS [56], Geo-NeuS [15], and NeuralAngelo [33]. We also compared it with recently emerged reconstruction methods based on 3DGS, such as SuGaR [19], 2DGS [21], and GOF [69]. All results are summarized in Fig. 5, Fig. 7, Fig. 8, Table II and Table III.
The DTU dataset: Our method achieves the highest reconstruction accuracy with relatively fast training speed. PGSR(DS) denotes downsampling to half the original image size for training. Our method significantly outperforms other 3DGS-based reconstruction methods. As shown in Fig. 7, our surfaces are smoother and contain more details.
@@
Table IV: Ablation study on the Meetingroom of TnT dataset.
{tabu}
@@
w/o Single-view 0.26 27.46
w/o Multi-view 0.15 28.14
w/o Our unbiased depth 0.20 26.80
Full model 0.29 27.30
@@ V-C Ablations
Our Unbiased Depth: From Fig 10, it can be observed that our overall geometric structure appears smoother and more precise, especially in flat regions. Table V-B also demonstrates that our depth rendering method achieves higher reconstruction and rendering accuracy.
Single-View and Multi-view Regularization: The single-view regularization term can provide a good initial geometric accuracy without relying on multi-view information. When single-view regularization is removed, the reconstruction accuracy decreases. Multi-view regularization effectively constrains the consistency of geometry between multiple views, improving overall reconstruction accuracy. From Table V-B, it is evident that multi-view regularization is crucial for reconstruction accuracy.
Table V: Ablation study on exposure Compensation.
{tabu}
[…]
w/o exposure modeling 0.76 21.71
w exposure modeling 0.80 25.77
Exposure Compensation: We validated the exposure compensation on the Ignatius series of the TnT dataset. As shown in Table V-C, exposure compensation enhances reconstruction and rendering quality.
@@ VI Limitations and Future Work
Although our PGSR efficiently and faithfully performs geometric reconstruction, it also faces several challenges. Firstly, we cannot perform geometric reconstruction in regions with missing or limited viewpoints, leading to incomplete or less accurate geometry. Exploring methods to improve reconstruction quality under insufficient constraints using priors is another avenue for further investigation. Secondly, our method does not consider scenarios involving reflective surfaces or mirrors, so reconstruction in these environments will pose challenges. Integrating with existing 3DGS work that accounts for reflective surfaces would enhance reconstruction accuracy in such scenarios. Finally, we found that there are some floating points in the scene, which affect the rendering and reconstruction quality. Integrating more advanced 3DGS baselines [39] would help further enhance overall quality.
@@ VII Conclusion
In this paper, we propose a novel unbiased depth rendering method based on 3DGS. With this method, we render the plane geometry parameters for each pixel, including normal, distance, and depth maps. We then incorporate single-view and multi-view geometric regularization, and exposure compensation model to achieve precise global consistency in geometry. We validate our rendering and reconstruction quality on the MipNeRF360, DTU, and TnT datasets. The experimental results indicate that our method achieves the highest geometric reconstruction accuracy and rendering quality compared to the current state-of-the-art methods.

### v2

@@ V-B Reconstruction
We compared our method, PGSR, with current state-of-the-art neural surface reconstruction methods including NeuS [57], Geo-NeuS [15], VolSDF [61], and NeuralAngelo [33]. We also compared it with recently emerged reconstruction methods based on 3DGS, such as SuGaR [19], 2DGS [21], and GOF [68]. All results are summarized in Fig. 5, Fig. 7, Fig. 8, Table II and Table III.
The DTU dataset: Our method achieves the highest reconstruction accuracy with relatively fast training speed. Our method significantly outperforms other 3DGS-based reconstruction methods. As shown in Fig. 7, our surfaces are smoother and contain more details.
@@
Table IV: Ablation study on the TnT dataset.
@@
w/o Single-View 0.49 27.02
w/o Multi-View 0.32 27.30
w/o Multi-View Geometric 0.49 27.07
w/o Multi-View Photometric 0.39 26.83
w/o Geometric Occlusion Estimation 0.28 21.70
w/o Our unbiased depth 0.38 26.47
w/o Exposure Compensation 0.49 25.33
Full model 0.52 26.73
@@ V-C Ablations
Our Unbiased Depth: From Fig 10, it can be observed that our overall geometric structure appears smoother and more precise, especially in flat regions. Table IV also demonstrates that our depth rendering method achieves higher reconstruction and rendering accuracy.
Single-View and Multi-View Regularization: The single-view regularization term can provide a good initial geometric accuracy without relying on multi-view information. When single-view regularization is removed, the reconstruction accuracy decreases. Multi-view regularization constrains the consistency of geometry between multiple views, improving overall reconstruction accuracy. Both multi-view photometric and geometric consistency contribute to improving reconstruction accuracy. From Table IV, it is evident that multi-view regularization is crucial for reconstruction accuracy. However, without incorporating potential occlusion estimation, the multi-view regularization term will have a negative effect, leading to poor surface reconstruction and rendering accuracy.
The ablation results also reflect another issue: geometric constraints slightly degrade rendering quality. We speculate that this is due to an incomplete image rendering model, which forces the system to strike a balance between image and geometry losses. Further exploration may be needed to achieve synchronized improvements in geometry and novel view synthesis.
Exposure Compensation: As shown in Table IV, exposure compensation enhances reconstruction and rendering quality.
@@ VI Limitations and Future Work
Although our PGSR efficiently and faithfully performs geometric reconstruction, it also faces several challenges. Firstly, we cannot perform geometric reconstruction in regions with missing or limited viewpoints, leading to incomplete or less accurate geometry. Exploring methods to improve reconstruction quality under insufficient constraints using priors is another avenue for further investigation. Secondly, our method does not consider scenarios involving reflective surfaces or mirrors, so reconstruction in these environments will pose challenges. Integrating with existing 3DGS work that accounts for reflective surfaces would enhance reconstruction accuracy in such scenarios. Finally, we found that there are some floating points in the scene, which affect the rendering and reconstruction quality. Integrating more advanced 3DGS baselines [39] would help further enhance the overall quality.
@@ VII Conclusion
In this paper, we propose a novel unbiased depth rendering method based on 3DGS. With this method, we render the plane geometry parameters for each pixel, including normal, distance, and depth maps. We then incorporate single-view and multi-view geometric regularization, and exposure compensation model to achieve precise global consistency in geometry. We validate our rendering and reconstruction quality on the MipNeRF360, DTU, and TnT datasets. The experimental results show that our method achieves the highest geometric reconstruction accuracy and competitive rendering quality compared to state-of-the-art methods.
Supplementary Document
I Additional Implement Details
I-A Details of the Multi-View Regularization
Building the image graph: We use the camera poses from the training data to calculate the relative angles and positions between all training frames, sorting the frame distances first by relative angle and then by position distance. Based on the specific characteristics of the dataset, we set the corresponding maximum number of neighboring frames, maximum relative angle, minimum relative position, and maximum relative position. For any given reference frame in the training data, other training frames that meet the set conditions are grouped into the neighboring frame set of that reference frame. For the DTU [23], TnT [28], and Mip360 [2] datasets, we uniformly set a maximum of 8 neighboring frames, a maximum relative angle of 30 degrees, a minimum relative position of 0.01, and a maximum relative position of 1.5.
Select Neighboring Frames During Training: During each training iteration, we randomly select one frame from the neighboring frame set of the reference frame to be used as the neighboring frame.
I-B Depth Filtering
There are often erroneous noise points near the edges of the depth map. We filter out these noise points using the angle between the rays and the normals estimated from the depth map:
$\theta=acos(\frac{\mid\bm{N}_{d}(\bm{p})\cdot\bm{V}(\bm{p})\mid}{\left|\bm{N}_{d}(\bm{p})\right|\left|\bm{V}(\bm{p})\right|}),$ (I1)
$\bm{N}_{d}(\bm{p})=\frac{(\bm{P}_{1}-\bm{P}_{0})\times(\bm{P}_{3}-\bm{P}_{2})}{|(\bm{P}_{1}-\bm{P}_{0})\times(\bm{P}_{3}-\bm{P}_{2})|},$ (I2)
For a pixel point $\bm{p}$, we sample four points from the neighboring pixels and project sampled depth points into 3D points $\{\bm{P}_{j}|j=1,...,4\}$ in the camera coordinate system, then calculate the normal $\bm{N}_{d}(\bm{p})$ for the pixel point $\bm{p}$. $\bm{V}(\bm{p})$ is the ray.
When $\theta$ is greater than $80^{\circ}$, we consider the depth point to be noise. After completing the training for each scene, we apply depth filtering as post-processing to the rendered depth maps and then perform surface reconstruction using the TSDF algorithm. Fig I5 shows that noise at the edges can be filtered out. We only applied depth filtering on the TnT dataset, and according to the ablation experiment results, normal-based depth filtering slightly improves surface reconstruction accuracy.
Figure I1: Qualitative comparison of surface normals under different single-view regularization terms. The baseline model does not include any geometric regularization terms. Adding the Single-View constraint without Image Edge-Aware loss on top of the baseline can produce relatively smooth surfaces, but some areas remain insufficiently flat. Incorporating the Image Edge-Aware Single-View constraint helps retain more details.
Figure I2: Qualitative comparison of different geometric regularization terms. Although the single-view regularization term can produce relatively smooth surfaces, it tends to create holes on highly specular metallic objects. Building on this, introducing a multi-view geometric regularization term, which enforces multi-view geometric consistency, can further improve surface smoothness, but some holes may still remain. By incorporating the complete geometric regularization term, it is possible to generate surfaces that are both smooth and rich in detail.
Figure I3: Qualitative comparison of multi-view photometric without and with NCC. Multi-view photometric consistency with NCC can reconstruct higher-quality surfaces.
Figure I4: Qualitative comparison of rendered image without and with exposure compensation. Exposure compensation can effectively adjust for overall brightness changes in the image. After compensation, the brightness of the adjusted image is closer to the ground truth image.
Figure I5: Qualitative comparison of point clouds without and with depth filtering. Depth filtering can effectively remove noise at the edges.
[…]
Barn Caterpillar Courthouse Ignatius Meetingroom Truck Mean
w/o Single-View 0.64 0.41 0.18 0.80 0.30 0.63 0.49
w/o Edge-Aware 0.65 0.44 0.20 0.81 0.32 0.65 0.51
w/o Multi-View 0.45 0.22 0.13 0.53 0.19 0.4 0.32
w/o Multi-View Geometric 0.65 0.41 0.16 0.8 0.32 0.61 0.49
w/o Multi-View Photometric 0.55 0.28 0.16 0.60 0.21 0.54 0.39
w/o Geometric Occlusion Estimation 0.48 0.03 0.17 0.44 0.18 0.40 0.28
w/o Our Unbiased Depth 0.5 0.31 0.18 0.70 0.15 0.45 0.38
w/o Exposure Compensation 0.66 0.42 0.20 0.77 0.29 0.62 0.49
w/o Depth Filter 0.66 0.43 0.20 0.80 0.33 0.60 0.50
Full model 0.66 0.44 0.20 0.81 0.33 0.66 0.52
[…]
Barn Caterpillar Courthouse Ignatius Meetingroom Truck Mean
w/o Single-View 29.45 27.06 23.89 26.68 28.14 26.9 27.02
w/o Edge-Aware 29.14 26.80 23.43 26.49 27.67 26.58 26.69
w/o Multi-View 29.82 27.11 24.3 26.89 28.51 27.14 27.30
w/o Multi-View Geometric 29.55 27.05 24.14 26.63 28.16 26.90 27.07
w/o Multi-View Photometric 29.33 26.91 23.62 26.61 27.75 26.76 26.83
w/o Geometric Occlusion Estimation 22.44 21.20 19.95 21.95 22.52 22.51 21.70
w/o Our Unbiased Depth 28.60 26.63 22.88 26.49 27.63 26.58 26.47
w/o Exposure Compensation 29.21 24.27 22.47 23.17 26.19 26.67 25.33
Full model 29.2 26.81 23.48 26.48 27.72 26.66 26.73
[…]
Barn Caterpillar Courthouse Ignatius Meetingroom Truck Mean
2DGS [21] 0.45 0.24 0.13 0.50 0.18 0.43 0.32
2DGS [21] + Multi-View 0.51 0.33 0.10 0.67 0.24 0.61 0.41
2DGS with expected depth [21] + Multi-View 0.64 0.39 0.13 0.75 0.31 0.61 0.47
II Additional Results
II-A Additional Detailed Ablation Experiments
Table I1 shows the ablation results for each scene in the TnT dataset. It can be seen that each modification is necessary and contributes to improving surface reconstruction accuracy.
Image Edge-Aware Single-View: From Table I1 and I2, it can be observed that single-view constraints help improve surface reconstruction accuracy but slightly reduce rendering quality. Edge awareness offers only a slight improvement in surface reconstruction, but as seen in the comparison between (c) and (b) in Fig I1, edge awareness is beneficial for preserving more details.
Multi-View Geometric Consistency: The qualitative effect of multi-view geometric consistency is shown in Fig I2. By enforcing multi-view geometric consistency, surface smoothness is improved. For highly specular metallic objects, it is difficult to ensure geometric consistency relying on images and single-view constraints. By imposing multi-view geometric consistency, the reconstructed surface becomes more complete and smooth. Table I1 also shows that multi-view geometric consistency improves surface reconstruction accuracy across various scenes, though it slightly reduces rendering quality, as demonstrated in Table I2.
Multi-View Photometric Consistency: The qualitative results of multi-view photometric consistency are shown in Figs I2 and I3. Fig I2 demonstrates that the proposed multi-view photometric consistency significantly improves surface reconstruction quality, producing smooth and detail-rich surfaces, especially in highly specular scenes. Fig I3 illustrates the benefits of using NCC (Normalized Cross-Correlation) to compute photometric consistency. In highly specular scenes, directly measuring photometric consistency with the absolute difference between two pixel patches is sensitive to pixel intensity variations, resulting in poor surface reconstruction quality. In contrast, NCC calculates similarity based on correlation coefficients, making it less sensitive to brightness changes and more suitable for image matching, thereby achieving better surface quality. Table I1 and I2 shows that in various scenes, the multi-view photometric consistency constraint plays a crucial role in improving reconstruction quality, though it slightly reduces rendering quality.
Unbiased Depth: Unbiased depth rendering is crucial for producing high-quality surface reconstructions. As shown in Table I1, unbiased depth rendering is essential for improving surface reconstruction quality across all scenes. Table I2 also demonstrates that unbiased depth rendering can slightly enhance rendering quality.
Exposure Compensation: We assign two exposure coefficients to each frame to model the exposure variation across frames. The qualitative results of the Ignatius sequence from the Tanks and Temples dataset are shown in Fig I4. Most frames of the Ignatius sequence are relatively bright, with a few being darker. Without exposure compensation, the rendered images tend to appear overly bright, whereas the images with exposure compensation have overall brightness closer to the ground truth. As seen from Table I1 and Table I2, exposure compensation helps improve both surface reconstruction and rendering quality.
Depth Filter: For the Tanks and Temples dataset, after training each scene, depth maps for each frame are rendered. Depth filtering is then applied to remove depth noise, followed by the use of the TSDF algorithm to fuse the cleaned depth maps and generate an SDF field, from which the mesh is extracted. Table I1 shows that depth filtering can slightly improve surface reconstruction quality. The ablation study Fig I5 demonstrates the effectiveness of depth filtering.
Related to 2DGS: For TnT scene, 2DGS [21] defaults to using median depth. 2DGS with expected depth, on the other hand, employs the expected depth mode. We conducted ablation comparisons between 2DGS and our multi-view regularization. As shown in the Table I3, our multi-view regularization applies to 2DGS, significantly enhancing the surface reconstruction accuracy of 2DGS. However, due to the ’disk-aliasing’ problem associated with median depth, the accuracy improvement is relatively limited compared to expected depth. Even with the addition of multi-view constraints, the surface accuracy of 2DGS is still inferior to that of PGSR.
II-B More Results
Figs II6 to II9 present a comparison of surface reconstruction results across various scenes between PGSR, 2DGS [21], and GOF [68]. We provide additional results on various scenes and objects, further confirming the capability of our method, PGSR, in achieving high-fidelity geometric reconstruction, as shown in Fig II10.
Figure II6: Qualitative comparisons in surface reconstruction between PGSR, 2DGS [21], and GOF [68] on the DTU dataset.
Figure II7: Qualitative comparisons in surface reconstruction between PGSR, 2DGS, and GOF on the DTU dataset.
Figure II8: Qualitative comparisons in surface reconstruction between PGSR, 2DGS, and GOF on the DTU dataset.
Figure II9: Qualitative comparisons in surface reconstruction between PGSR, 2DGS, and GOF.
Figure II10: PGSR achieves high-precision geometric reconstruction in various indoor and outdoor scenes from a series of RGB images without requiring any prior knowledge.